Hi all, I was hoping to get some help with the following problem.
Find ∆y/∆x (given the following information)
b) y = sqrt(x), x=9, ∆x= 0.9
Any help would be great.
-Nick.

Answer by HyperBrain(694)

Sorry, I don't know how to input the delta sign.
x=9 so y=sqrt(9)=3
since x=9, delta_x= 0.9, then x + delta_x=9.9
y + delta_y=sqrt (x + delta_x)= sqrt (9.9)=3.146 (approx.)
---
delta_y=(y + delta_y)-y=3.146-3=0.146
---
so, delta_y/delta_x=0.146/0.9=0.0162
---